OpenStudy (anonymous):

simplify tan^4x+2tan^2 + 1

OpenStudy (anonymous):

Is this tan x to the power 4 plus 2 tanx squared plus 1?

OpenStudy (anonymous):

hint: factorise and apply pythagorean identities.

OpenStudy (anonymous):

\[\tan^2 x +1 = \sec^2 x\] \[x^2 + 2x +1 = (x+1)^2\] I believe these hints are sufficient. Good luck!